OpenStudy (anonymous):

2 people walk 4 miles toward each other, one at 3mph and the other at 4mph. Show an equation that shows how long it will take before they meet.

OpenStudy (anonymous):

Say they meet at time t hours when one will have walked 3t miles and the other 4t miles so 7t = 4 and t= 4/7 hours

OpenStudy (anonymous):

how do u translate that to minutes and show what time they would meet

OpenStudy (anonymous):

They meet after 4/7 hours. There are 60 minutes in an hour last time I checked so multiply the hours by 60.

OpenStudy (anonymous):

can u show me?

OpenStudy (anonymous):

No.

OpenStudy (anonymous):

I got the answer t= 4/7 but I am having a difficult time converting 3t and 4t into minutes, won't you help?

OpenStudy (anonymous):

Why u want to do that?

OpenStudy (anonymous):

I tried using 240 minutes and tried 57600 seconds. To find out what time they meet.

OpenStudy (anonymous):

I keep telling u they meet after 4/7 hours (ie a bit over half hour after they start walking)

OpenStudy (anonymous):

I know but I need a specific time. Thanks for your patience

OpenStudy (anonymous):

4/7 hours, how many more times!!!!!

OpenStudy (anonymous):

What do u mean by specific, 4/ hours is VERY specific.

OpenStudy (anonymous):

4/7

OpenStudy (anonymous):

Well?

OpenStudy (anonymous):

Before next week would be nice...

OpenStudy (anonymous):

the equation is 3t+4t =4 but they meet after one has gone 12/7 and the other has gone 16/7 so i don't see how solving t gives me a specific time

OpenStudy (anonymous):

3 miles per hour for t hours = distance in miles for one of them 4 miles per hour for t hours = distance in miles for the other They walk in total 4 miles so 7t = 4

OpenStudy (anonymous):

Anything else?

OpenStudy (anonymous):

I got that but what time would they meet. 12/7 of 4 hours?

OpenStudy (anonymous):

Duh... if 7t = 4 t= 4/7 hours for the umpteenth time.

OpenStudy (anonymous):

that's as far as we can go, Thanks